OpenStudy (anonymous):

Write the equation y - 2 = 2(x - 3) in standard form. Answer y - 2x = 4 2x - y = -6 2x + y = 6 2x - y = 4

OpenStudy (anonymous):

@hba by standard form do they mean y=mx+b?

OpenStudy (hba):

Standard form is, \[\huge\ Ax+By+C=0\]

OpenStudy (hba):

Where C is the constant :)

OpenStudy (anonymous):

?

OpenStudy (hba):

Um arrange it in this order. 1) expand x terms out and get all of them on Left hand side

OpenStudy (anonymous):

yeah im comply lost dude

OpenStudy (hba):

Where are you confused ?

OpenStudy (anonymous):

the whole thing like as soon as ya put up the standard form it lost me cuz where r we getting the a b and c?

OpenStudy (hba):



OpenStudy (anonymous):

so y-2=2x-6?

OpenStudy (hba):

Let me show you an example 6x+5y-6=0 ^ like this So its standard form we write Ax+By+C=0

OpenStudy (hba):

y-2=2x-6 Add 6 on both sides :)

OpenStudy (anonymous):

y-8=2x?

OpenStudy (hba):

-2+6=

OpenStudy (anonymous):

oh 4 so y+4=2x?

OpenStudy (hba):

Now subtract (y+4) from both sides :)

OpenStudy (anonymous):

wait how do i do that if their is only 2x?

OpenStudy (hba):

(y+4)-(y+4)=2x-(y+4)

OpenStudy (anonymous):

???? so it would give me 2x-y+4

OpenStudy (hba):

0=2x-y-4 or 2x-y-4=0 ^ Answer

OpenStudy (anonymous):

none of my answers at the top match that though

OpenStudy (anonymous):

@hba

OpenStudy (hba):

Add both sides by 4 then :)

OpenStudy (anonymous):

oh so it would be 2x-y=4

OpenStudy (phi):

here are various forms of the equation of a line http://en.wikipedia.org/wiki/Linear_equation#General_.28or_standard.29_form